# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ 2.1 Models and Estimations
$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (2.2)
@@
where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$. For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^3] We use a model with a linear index solely for the purpose of illustration. Our method applies to general models. $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
@@
where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$. In addition, when $X_{1,it}=1$, the DGP in (2.2) consists of the traditional individual- and time-effects in the intercept. When $X_{k,it}=Y_{it-1}$ for some $k$, the DGP is dynamic because each $Y_{it}$ depends on its past value.
@@
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (2.3)
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (2.4)
@@
$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (2.5)
for some $3\leq K_{1}<K$. To our knowledge, except (2.3) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors. The specification in (2.4) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model. Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (2.4) only for strictly exogenous $X_{k,it}$. Currently, no methods are available for nonlinear models with (2.5).
We are interested in the estimation of and the inference about $\theta_{0}$ in (2.1), under the presence of the nuisance parameter $\phi^{0}$. Our approach is likelihood-based. Denote $\phi:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$ and define the log-likelihood function (“likelihood” hereafter) as
where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}$. This rules out the indeterminancy in (2.7) without explicitly using the Lagrangian. We use the following example to illustrate the reparameterization.
Example 2 (reparameterization)
For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$. In Example 3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
@@
$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (2.8)
Notice that (2.8) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (2.8) over $i=1,\ldots,N-1$, we obtain
$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (2.9)
Combining (2.8) and (2.9), it follows that
@@
In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$. Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively. For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
@@
Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (2.9), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
2.2 Incidental Parameter Problem
The dynamic nonlinear heterogeneous parameter, DN-HP, models generally suffer from the incidental parameter problem (IPP). We now discuss this problem. We use $\mathbb{E}$ to denote the expectation w.r.t.
$\prod_{i=1}^{N}\prod_{t=1}^{T}f(y_{it}|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$
which is the true density evaluated at $\theta_{0}$, given: i) initial values of the predetermined regressors, ii) all strictly exogenous regressors, and iii) the unobserved effect $\phi^{0}$.
Remark 1 (expectation)
Following the profiled likelihood framework of Pace and Salvan (2006), we continue our discussion by defining, for a given $\theta$ and each pair of $(N,T)$,
are the score and Hessian. The remainder term $r^{l}(\theta)$ is characterized by Equation (A.17) in Appendix A. Here the estimation errors $\widehat{\psi}(\theta)-\psi(\theta)$ can be derived by another Taylor expansion. Seeing $s(\theta,\widehat{\psi}(\theta))=0$ as $\widehat{\psi}(\theta)$ is the maximizer, a Taylor expansion of $s(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$ gives
00 0 $=s\left(\theta\right)+H\left(\theta\right)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{s}(\theta),$
where $r^{s}(\theta)$ is the remainder term of the expansion, which is given by Equation (A.6) in Appendix A. Combining (2.13) and (2.12), and taking expectation, we obtain
$\sqrt{NT}\mathbb{E}[l(\theta,\widehat{\psi}(\theta))-l(\theta,\psi(\theta))]=-\sqrt{NT}B(\theta)+\sqrt{NT}\mathbb{E}r(\theta),\qquad B(\theta):=\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\},$ (2.14)
where $r(\theta)$ is a remainder term depending on $r^{l}(\theta)$ and $r^{s}(\theta)$, given by Equation (A.16) in Appendix A. By Lemma A.3 in Appendix A, we show that $\mathbb{E}r(\theta)$ is of order $o(1/\sqrt{NT})$, uniformly in $\theta$, as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$. This implies that $\sqrt{NT}\mathbb{E}r(\theta)=o(1)$ and is negligible. $B(\theta)$ is the IPP bias arising from the estimation errors of $\widehat{\psi}(\theta)$. As opposite to $\mathbb{E}r(\theta)$, the bias $B(\theta)$ is not negligible, in the sense that $\sqrt{NT}\mathbb{E}B(\theta)\not\rightarrow 0$ under the above asymptotic sequences $\left\{N,T\right\}$ (see Theorem 1). Consequently, $\widehat{\theta}$ and many likelihood-based test statistics inherit the bias from the likelihood, leading to invalid inferences. We consider the asymptotic sequences $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$, because $\sqrt{NT}B(\theta)$ explodes under $N/T\rightarrow 0$ or $N/T\rightarrow\infty$. We illustrate the IPP and its consequence to the structural parameter estimator in the example below, together with the issue of $N/T\rightarrow\kappa$.
Example 3 (heterogeneous mean model)
Following the same setting in Example 2 but let $\varepsilon_{it}$ be normal with mean 00 and variance $\sigma_{0}>0$. In the current example, $\theta_{0}=(\mu_{0},\sigma_{0})^{\prime}$ is the structural parameter of interests. This example is a modified version of Neyman and Scott (1948) and is similar to Fernández-Val and Weidner (2016), except that the mean of $Y_{it}$ is $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, conditional on $\phi^{0}$.
Define [^4] In this simple example, $\varepsilon_{it}$ is independent across both $i$ and $t$. Therefore, $\mathbb{E}$ is simply w.r.t. the normal density having mean $\mu^{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and variance $\sigma^{0}$. $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}$. Following the same steps in Example 2, we can derive
@@ 2.2 Incidental Parameter Problem
$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (2.15)
@@
$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (2.16)
@@
$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (2.17)
where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (2.15) and (2.16) into (2.17), we obtain, after some algebra,
@@
$l(\theta,\widehat{\psi})$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log\sigma-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2},$
$\partial_{\sigma}l(\theta,\widehat{\psi})$ $=-\frac{1}{2\sigma}+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma^{2}}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}.$
Following the same steps in Example 2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (2.10); and it is independent of $\sigma$. Therefore, equating $\partial_{\sigma}l(\theta,\widehat{\psi})=0$, substituting $\mu$ with $\widehat{\mu}$, and solving for $\sigma$, the estimator of $\sigma^{0}$ is

### v5

@@ 5 Empirical Analysis
$Y_{it}$ $=\mathds{1}\!\left\{(\rho+\zeta_{i}+\eta_{t})Y_{i,t-1}+(\beta+\alpha_{i}+\gamma_{t})X_{it}+c_{i}+\tau_{t}+\varepsilon_{it}>0\right\},$ (5.2)
where $\varepsilon_{it}$ follows the standard normal distribution, and $\sum_{i}c_{i}=\sum_{i}\zeta_{i}=\sum_{i}\alpha_{i}=\sum_{t}\eta_{t}=\sum_{t}\gamma_{t}=0$. Model (5.1) is the dynamic homogeneous-slope model, which includes two-way fixed effects in the intercept. Model (5.2) is the dynamic heterogeneous-slope model, which extends the specification by allowing the slope coefficients—that is, the coefficients on both the lagged dependent variable and the number of children—to vary across individuals and over time: $\rho_{it}=\rho+\zeta_{i}+\eta_{t}$ and $\beta_{it}=\beta+\alpha_{i}+\gamma_{t}$. The parameter $\beta_{it}$ measures the heterogeneous impact of the number of children on a mother’s latent propensity to work, while $\rho_{it}$ captures heterogeneous state dependence in labor force participation. Model (5.2) nests Model (5.1) as a special case when $\rho_{it}$ and $\beta_{it}$ are constant across $i$ and $t$.
As discussed in Section 1, when the true effect of $X_{it}$ on the outcome, $\beta_{it}$, varies across the population, it is important to account for its potential correlation with $X_{it}$. In practice, such correlation often arises from self-selection. For instance, single mothers who choose to have more children might be those with greater financial resources or better access to childcare, such that their labor supply is less affected by additional children. Alternatively, correlation between $\beta_{it}$ and $X_{it}$ may stem from common time trends: over time, fertility rates declined while concurrent factors such as rising childcare and schooling costs, advances in home technology, and evolving welfare policies (e.g., child tax credits) altered the effect of each additional child on their mother’s LFP. In both scenarios, Model (5.2) is the appropriate specification. It controls for unobserved heterogeneity in $\beta_{it}$ by incorporating individual and time fixed effects into the slope and allowing them to correlate with $X_{it}$. Controlling for heterogeneous state dependence further improves robustness by allowing the persistence of labor-force participation to differ across individuals and periods, capturing additional sources of dynamic heterogeneity.
Given our small sample size, direct estimation of both models could suffer from severe incidental parameter problems, which we address using our bias correction procedure. Table 3 presents the estimation results. Panel A reports results for Model (5.1) under the homogeneous slope assumption. For each parameter, we provide the MLE and bias-corrected estimates, along with their standard errors. Examining the bias-corrected estimates, we observe that an increase in the number of children appears to reduce a single mother’s likelihood of labor force participation, while past LFP status strongly predicts current LFP. However, this relationship shifts dramatically once we account for slope heterogeneity. Panel B presents the results for Model (5.2), allowing for heterogeneous slopes. Here, the bias-corrected estimates reveal a positive average impact of additional children on a mother’s labor supply: each additional child increases the latent propensity to participate in the labor force by 0.436, which corresponds to an average increase of about 7 percent in the probability of employment. This indicates that, on average, single mothers with more children are more likely to work, likely driven by the increased financial demands of supporting multiple children independently. Additionally, unlike the homogeneous-slope model, the serial correlation in labor-force participation remains statistically significant but diminishes in magnitude, suggesting that what appears to be strong state dependence in the homogeneous model partly reflects unobserved heterogeneity in individuals’ persistence of employment rather than uniform structural dynamics. Together, these findings underscore the importance of controlling for unobserved slope heterogeneity, as failing to do so can lead to substantially different conclusions.
In addition to highlighting the difference between homogeneous and heterogeneous slope models, Table 3 demonstrates the importance of bias-correction in estimating dynamic nonlinear panel data models. Examining the results in Panel B, we observe that the MLE overestimates the impact of the number of children while underestimating the degree of serial correlation in LFP. Bias correction results in a 38 percent lower estimate of the former and a 32 percent higher estimate of the latter.
Finally, in Figure 2(a), we plot the distribution of the estimated $\beta_{it}$ from Model (5.2), with summary statistics in Table G.2 (supplementary appendix). Each $\beta_{it}$ estimate is constructed as the sum of the bias-corrected average effect $\beta$ and the estimated individual and time effects $\alpha_{i}$ and $\gamma_{t}$. These estimates reveal substantial heterogeneity in how the number of children influences single mothers’ labor force participation. While the mean effect is positive, about 46 percent of the individual-time observations show a decrease in LFP with each additional child, suggesting that the average effect alone inadequately captures the diverse impacts across individuals and over time. Figure 2(b) illustrates the temporal evolution of these heterogeneous effects. The median $\beta_{it}$ rises steadily from 1987 to 1994—implying an increasingly positive labor-supply response to additional children—before declining modestly thereafter. The interquartile and decile bands remain wide throughout, highlighting persistent dispersion across individuals. Together, these results reinforce the importance of accounting for both individual and temporal variations in slope coefficients, as the impact of children on labor force participation can vary significantly based on unobserved factors unique to each mother and time period.
   Number of Children    Lagged Participation
Uncorrected Corrected Uncorrected Corrected
      Panel A. Homogeneous Slope
Model parameter -0.923 -0.851 0.589 0.540
(0.232) (0.228) (0.108) (0.107)
      Panel B. Heterogeneous Slope
Model parameter 0.704 0.436 0.191 0.253
(0.141) (0.144) (0.106) (0.105)
  • Notes: Standard errors in parentheses. Panel A reports the estimation results of the dynamic homogeneous-slope model. Panel B reports the estimation results of the dynamic heterogeneous-slope model. The “uncorrected” and the “corrected” columns show respectively the MLE and the bias-corrected parameter estimates. Data source: PSID 1987–1997.
Figure 2: Estimated Heterogeneous Effects of Children on Single Mother Labor Force Participation
(a) Distribution of Estimated Effects
(b) Temporal Evolution of Estimated Effects
[…]
Notes: Panel (a) shows the distribution of the estimated individual-time effects $\beta_{it}$ from the dynamic heterogeneous-slope model. Panel (b) plots the median (solid), interquartile range (25th–75th, dashed), and decile range (10th–90th, dotted) of $\beta_{it}$ over time.
In conclusion, our analysis demonstrates the efficacy of our likelihood-based bias correction method for estimating dynamic nonlinear models of labor supply with multi-dimensional heterogeneities. We show the importance of controlling for unobserved slope heterogeneity when it may correlate with regressors and the value of panel data models that accommodate two-way fixed effects in both the intercept and the slope. Our findings reveal that the number of children has, on average, a positive influence on a single mother’s labor force participation, but such effect is highly heterogeneous across individuals and time periods. This heterogeneity underscores the complex nature of labor supply decisions and the unique challenges each single mother faces in balancing work and child-rearing responsibilities.
6 Final Remarks
In this paper, we propose a likelihood-based analytical bias correction procedure for a general class of two-way dynamic nonlinear heterogeneous parameter models subject to the incidental parameter problem. We give the analytical form of a corrected likelihood and show that it delivers point estimators that are asymptotically unbiased and test statistics that are asymptotically $\chi^{2}$-distributed. Simulation studies and empirical analyses support our claims.
Several issues deserve further studies. First, we have not rigorously investigated the asymptotic properties of the average partial effects. Admittedly, average partial effect is very important, especially for nonlinear models. It is therefore worthwhile to formally establish its asymptotic theory. Second, there has been various fixed-$T$ consistent solutions for the logit model with fixed effects. However, to the best of our knowledge, we have not seen any such developments for the two-way heterogeneous parameter logit model. It could be useful to push such a development, because the logit model is widely used in many studies. Finally, certain micro panels tend to have a small $T$, in which case the higher-order bias terms may be non-negligible. For this reason, it may be potentially important to develop high-order bias correction approaches for two-way dynamic nonlinear heterogeneous parameter models.
Acknowledgement
We would like to thank the editor Iván Fernández-Val and two anonymous referees for their valuable comments and suggestions. We would like to express our gratitude to comments and suggestions from Geert Dhaene, Yanqin Fan, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu. Our research assistants Yuchi Liu and Xinrui Zhang helped us with various tasks. Leng’s research is partially supported by the National Natural Science Foundation of China (72573136) and the Fujian Natural Science Foundation in China (2024J08013). Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
Appendix
[…]
AppendixAppendix
A Examples
Example A.1 (heterogeneous parameter model)
Consider a probit model with the data-generating process
$Y_{it}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (A.1)
@@ A Examples
where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$. For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^5] We use a model with a linear index solely for the purpose of illustration. Our method applies to general models. One example is the modified Neyman-Scott model in Example A.3. $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
@@
where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$. In addition, when $X_{1,it}=1$, the data-generating process in (A.1) consists of the traditional individual- and time-effects in the intercept. When $X_{k,it}=Y_{it-1}$ for some $k$, the data-generating process is dynamic because each $Y_{it}$ depends on its past value.
@@
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (A.2)
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (A.3)
@@
$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (A.4)
for some $3\leq K_{1}<K$. To our knowledge, except (A.2 ) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors. The specification in (A.3) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model. Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (A.3) only for strictly exogenous $X_{k,it}$. Currently, no methods are available for nonlinear models with (A.4).
Example A.2 (reparameterization)
For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$. In Example A.3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
@@
$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (A.5)
Notice that (A.5) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (A.5) over $i=1,\ldots,N-1$, we obtain
$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (A.6)
Combining (A.5) and (A.6), it follows that
@@
In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$. Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange-multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively. For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
@@
Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (A.6), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
Example A.3 (heterogeneous mean model)
Following the same setting in Example A.2 but let $\varepsilon_{it}$ be normal with mean 0 and variance $\sigma_{0}>0$. In the current example, $\theta_{0}=(\mu_{0},\sigma_{0})^{\prime}$ is the structural parameter of interests. This example is a modified version of Neyman and Scott (1948) and is similar to Fernández-Val and Weidner (2016), except that the mean of $Y_{it}$ is $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, conditional on $\phi^{0}$.
Define [^6] In this simple example, $\varepsilon_{it}$ is independent across both $i$ and $t$. Therefore, $\mathbb{E}$ is simply w.r.t. the normal density having mean $\mu^{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and variance $\sigma^{0}$. $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}$. Following the same steps in Example A.2, we can derive
@@
$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (A.8)
@@
$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (A.9)
@@
$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (A.10)
where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (A.8) and (A.9) into (A.10), we obtain, after some algebra,
@@
[…]
Following the same steps in Example A.2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (A.7); and it is independent of $\sigma$. Therefore, equating $\partial_{\sigma}l(\theta,\widehat{\psi})=0$, substituting $\mu$ with $\widehat{\mu}$, and solving for $\sigma$, the estimator of $\sigma^{0}$ is